OpenStudy (anonymous):

) Question 9 Saved Solve for x: -2(x-1)=-6

OpenStudy (debbieg):

Distribute the -2 through the parentheses on the left side. What do you get?

OpenStudy (debbieg):

Then you will need to isolate the x. So use the addition/subtraction properties to move any non-x terms to the right side, then divide both sides by the coefficient of x.

OpenStudy (anonymous):

\[-2(x-1)=-6 \rightarrow (x-1)=\frac{-6}{-2} \rightarrow (x-1)=3 \] x= 3+1 x=4 is the required solution

OpenStudy (debbieg):

I could have told him that. Thought he might learn more by working through the solution on his own. Crazy thought.